# Indirect forces in disc-planet interaction

### v1

@@ 4.3 Deposition torque density
The decay of AMF $F_{J}$ relative to integrated torque $T$ in Figure 7 at large $R$ is caused by the dissipation of the density wave. Our simulations do not have any explicit sources of wave dissipation (e.g. viscosity or radiation), and numerical dissipation is very small, so that wave damping is driven predominantly by nonlinear effects — steepening of the wave profile leading to its shocking and dissipation (GR01; R02). This mechanism is effective even for low $M_{\mathrm{p}}$, when the nonlinearity is weak. But in a purely linear calculation this process and the associated wave dissipation are absent by default, as Figure 6 demonstrates.
@@
as follows from the wave angular momentum conservation. Non-zero $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ is what ultimately drives disc evolution, leading to phenomena such as gap formation (R02b; Kanagawa2015; Cordwell2024). Thus, knowing $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ is crucial for understanding evolution of planet-induced disc structures. This motivates us to examine the effect of the IT on the deposition torque density.
@@
We also show $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ obtained in a very different way, namely in a simulation with the same parameters as in Figure 7 but carried out in the barycentric frame using Disco, see Section 2.3. In this frame indirect forces are absent by default, meaning that this calculation provides an independent check of our astrocentric simulations. The result is shown by the red dot-dashed curve (as a function of barycentric distance, which should not make much difference), and matches the overall shape of $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ quite well. There are some differences, e.g. $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ in the barycentric frame is slightly larger than in the astrocentric frame, but nevertheless the agreement is quite good given the significant differences between the numerical methods used in this comparison.
@@ 4.4 Oscillatory behaviour of the indirect torque
where $H_{\mathrm{p}}$ is the disc scale height at the planetary location. For a disc illustrated in Figure 4, $q=1$ and the distance between the nodes decreases as $R_{k}^{-1/2}$, which is noticeable in the behavior of the green dot-dashed curve in panel (c) of that figure. Equations (25)-(26) are essentially analogous to the periodicity conditions formulated for the torque wiggles in disc-planet coupling (Cimerman2024b) and for the radial oscillations of the gravitational torque density in circumbinary discs (see Eqs. (17), (18) of Cimerman2024a) since all of these phenomena depend on the global spatial structure of the planet-driven density wave.
Furthermore, the dependence of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ on only the $m=1$ component of $\delta\Sigma$ allows us to explain the intriguing discrepancy between the linear and simulation-based amplitudes of the $\mathrm{d}T/\mathrm{d}R$ oscillations seen in Figure 5. Naively, one might expect the simulation-based $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations (solid) to have lower amplitude than the one predicted by the linear theory (dot-dashed), since in simulations the density wave amplitude decays as $R$ increases due to shock dissipation (see Figure 2a). However, Figure 2c shows that $\delta\Sigma_{1}$ actually increases with $R$, unlike all of the other $\delta\Sigma_{m}$ components. This phenomenon is caused by the nonlinear broadening of the wake as it propagates through the disc, which effectively transfers some power into the $m=1$ mode of the perturbation while the modes with $m>1$ get gradually extinguished by shock damping. This process has been previously described in Cimerman2024b and is illustrated by their Figure 14. As a result, far from the planet the simulation-based $\delta\Sigma_{1}$ exceeds the linear prediction for $\delta\Sigma_{1}$, explaining the trends in $\mathrm{d}T/\mathrm{d}R$ behavior observed in Figure 5.
@@ 4.5 Divergence of the indirect torque as $R\to\infty$
We now address the origin of the divergence of the indirect torque density $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ as $R\to\infty$, which is clearly visible in Figure 4. In principle, the reason for this divergence is already obvious from equation (14): according to its definition, $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ is proportional to $R^{2}$, which drives the divergence in the first place. However, the actual asymptotic behavior of $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ also depends on the behavior of $\delta\Sigma(R)$, which is something that we can address now.
Equation (24) shows that the amplitude of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations depends on $R$ as $\propto R^{2}|\delta\tilde{\Sigma}_{1}(R)|$, i.e. it increases with $R$ as long as $|\delta\tilde{\Sigma}_{1}(R)|$ decays slower than $R^{-2}$. This is clearly the case in the examples shown in Figures 4c, 6b, 7b, since the oscillation amplitude grows in a divergent fashion as $R\to\infty$. In the linear regime, in the absence of dissipation, one can show (R02; Miranda2020I) that far in the outer disc, where $\Omega\ll\Omega_{\mathrm{p}}$,
@@
After the wave shocks (which happens earlier for higher $M_{\mathrm{p}}$), the nonlinear wave dissipation causes $|\delta\tilde{\Sigma}_{1}(R)|$ to fall below the linear prediction (27). However, this has a rather limited effect in slowing down the growth of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations [^6] As a result of nonlinear damping, the wave AMF would eventually decay as $F_{J}\propto t^{1/2}$, where the function $t(R)$ is given by the equation (32) of R02. One can show that in this regime the amplitude of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations would scale as $\propto R^{(11/8)+(7q/16)-(5p/8)}$, not very far from the linear case (27).. Moreover, as Figure 2c illustrates, in low-$M_{\mathrm{p}}$ cases $|\delta\Sigma_{1}|$ may actually increase with $R$ (due to the transfer of power to $m=1$ mode as discussed in the end of Section 4.4) easily enabling the divergence of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations in Figure 4c.
On a more intuitive basis, one can show that the divergence of the indirect torque arises because of the reflex motion of the central star. Indeed, let us a consider a mass (e.g. a fluid element in the disc) in the barycentric frame of a star-planet system with a finite specific angular momentum ${\bf l}_{\mathrm{b}}={\bf r}\times\dot{\bf r}$, where $\bf r$} is the barycentric position of this mass. In the astrocentric coordinates the position of this mass is ${\bf R}={\bf r}-{\bf r}_{\star}$, where $\bf r$$_{\star}$ is the stellar position in the barycentric frame. Then the specific angular momentum ${\bf l}_{\mathrm{a}}$ of that mass in the astrocentric frame is
Moreover, because of planet-induced stellar motion around the barycenter of the system, $\dot{\bf r}_{\star}$ varies in time and so does ${\bf l}_{\mathrm{a}}$, with an infinitely growing amplitude as $r\to\infty$. This implies that in the astrocentric frame an infinitely large torque must be applied to that distant mass, thereby explaining the divergence of the indirect torque for $r,R\to\infty$ as a consequence of stellar reflex motion.
@@ 5 Torque on the planet and planetary orbital evolution
We start by emphasizing that in this Section the IT arises because of the reflex motion of the star caused by the gravity of the disc, and not of the planet as was the case in Sections 3 and 4. Taking a cross product of equation (2) with $M_{\mathrm{p}}{\bf R}_{\mathrm{p}}$ one finds that the torque on the planet due to the disk (its only non-zero $z$-component $T_{\mathrm{p}}$) is (Fairbairn2025a)
$T_{\mathrm{p}}=GM_{\mathrm{p}}\int\limits_{\rm disk}\Sigma({\bf R})\left({\bf R}_{\mathrm{p}}\times{\bf R}\right)_{z}\left(\frac{1}{|{\bf R}_{\mathrm{p}}-{\bf R}|^{3}}-\frac{1}{R^{3}}\right)\mathrm{d}^{2}{\bf R}\,$, (29)
@@
$T_{\mathrm{p,d}}=GM_{\mathrm{p}}\int\limits_{\rm disk}\Sigma({\bf R})\frac{\left({\bf R}_{\mathrm{p}}\times{\bf R}\right)_{z}}{|{\bf R}_{\mathrm{p}}-{\bf R}|^{3}}\,\mathrm{d}^{2}{\bf R}=-\int\limits_{\mathrm{disc}}\frac{\mathrm{d}T_{\mathrm{d}}(R^{\prime})}{\mathrm{d}R}\,\mathrm{d}R^{\prime}$ (30)
@@
$T_{\mathrm{p,id}}=-GM_{\mathrm{p}}\int\limits_{\rm disk}\Sigma({\bf R})\frac{\left({\bf R}_{\mathrm{p}}\times{\bf R}\right)_{z}}{R^{3}}\,\mathrm{d}^{2}{\bf R}.$ (31)
@@
Moreover, this also means that the full angular momentum of the disc-planet system is not conserved in the astrocentric frame. This result, which may seem paradoxical at first sight, is in fact fully in line with the findings of Rafikov2025, who showed that the angular momentum of a system of gravitating objects — the star, planet and disc, in our case — in a non-inertial frame attached to one of them, is not conserved because of the torque exerted on the system by the indirect force.
Until recently, studies computing the disc torque on the planet accounted only for the direct contribution $T_{\mathrm{p,d}}$, neglecting the indirect part $T_{\mathrm{p,id}}$ altogether. Fairbairn2025a explicitly considered $T_{\mathrm{p,id}}$ in their semi-analytical linear study of orbital evolution of low-mass eccentric planets and showed it to be small compared to $T_{\mathrm{p,d}}$. A similar conclusion was subsequently reached in Crida2025 based on simulations in the context of Type I migration, whereas for more massive (gap-opening) planets they found $T_{\mathrm{p,id}}$ to play a more significant role. We now re-assess the relative contribution of the indirect component $T_{\mathrm{p,id}}$ of the disc torque on the planet in our setup using (semi-)analytical linear calculations.
@@ 5.1 Low mass planets
To evaluate the relative contribution of $T_{\mathrm{p,id}}$ to $T_{\mathrm{p}}$, we start by recalling that the one-sided direct planetary torques on the disc are $\sim F_{J,0}$ and similar in magnitude but opposite in sign. The full direct torque given by equation (30) then ends up being (Ward1997)
@@
compare with equations (20) and (21). We note that for $R-R_{1}\gtrsim H_{\mathrm{p}}$, outside the $m=1$ OLR at $R_{1}$, $\Sigma_{1}=\delta\Sigma_{1}$ has a form given by the equations (22) and (23). As a result, the integrand in (36) rapidly oscillates in this region, so that this part of the disc provides a small contribution to $T_{\mathrm{p,id}}$. On the other hand, for for $R_{1}-R\gtrsim H_{\mathrm{p}}$, inwards of the $m=1$ OLR, $\delta\Sigma_{1}$ decays exponentially as $R$ decreases (see Figure 2b), except at $R=R_{\mathrm{p}}$ where the perturbation is co-orbital with the planet and becomes large. As a result, one expects that the main contributions to the integral in equation (36), as well as in equation (31), come from the immediate vicinity of the $m=1$ OLR, i.e. from $|R-R_{1}|\lesssim H_{\mathrm{p}}$, and from the corotation region $R\approx R_{\mathrm{p}}$. Denoting these contributions as $T_{\mathrm{L,id}}$ (Lindblad) and $T_{\mathrm{c,id}}$ (corotation), respectively, we can thus approximate $T_{\mathrm{p,id}}\approx T_{\mathrm{L,id}}+T_{\mathrm{c,id}}$. We corroborate these expectations in Appendix C, where we also provide an analytical calculation of $T_{\mathrm{L,id}}$.
@@
Fairbairn2025a studied the torque on the planet from the disc and the resultant planetary orbital evolution using semi-analytical linear theory (Miranda2019II; Fairbairn2022), as described in Section 2.4. We used their approach to numerically compute both $T_{\mathrm{p,id}}$ and $T_{\mathrm{p,d}}$ and plot their ratio in Figure 9 as a function of different disc parameters: temperature (a) and $\Sigma$ (b) slopes $q$ and $p$, respectively, and the disc aspect ratio $h_{\mathrm{p}}$, for $q=0$ and two values of $p=0$ (c) and $p=1.5$ (d).
@@
To shed light on these trends, we note that in a Keplerian disc with $p=1.5$, one expects the corotation torque to be absent, $T_{\mathrm{c,id}}\to 0$ (GT80). This is discussed in more detail in Appendix C, and is illustrated in Figure 11b. As a result, for this $p$ the indirect torque on the planet is provided only by the Lindblad contribution, $T_{\mathrm{p,id}}\approx T_{\mathrm{L,id}}$, which is known to be independent of $c_{\mathrm{s}}$ or $h_{\mathrm{p}}$ to zeroth order (see GT79; GT80), as underlined by the discussion in Appendix C and our analytical estimate (51). As a result, it follows from equation (35) that $|T_{\mathrm{p,id}}/T_{\mathrm{p,d}}|\propto h_{\mathrm{p}}^{2}$ for $p=1.5$, in agreement with Figure 9d.
@@
On the other hand, in a disc with $p=0$ the indirect corotation torque $T_{\mathrm{c,id}}$ is non-zero and significantly complicates the interpretation of the resultant trends (a similar conclusion has been reached by Crida2025). We find that in such a disc $T_{\mathrm{c,id}}$ provides a dominant contribution to $T_{\mathrm{p,id}}$ (e.g. see Figure 11a), strongly exceeding that of $T_{\mathrm{L,id}}$. For that reason, $T_{\mathrm{p,id}}$ has a higher amplitude for the $p=0$ disc as compared to the $p=1.5$ disc (the changing relative contributions of $T_{\mathrm{c,id}}$ and $T_{\mathrm{L,id}}$ cause the particular dependence on $p$ shown in Figure 9b). However, it has to be remembered that the magnitude of the corotation torque is a strong function of the potential softening, which itself depends on $h_{\mathrm{p}}$ (our linear calculations here adopt a softening length of $0.3h_{\mathrm{p}}$, following Fairbairn2025a); hence, there is some ambiguity in the actual amplitude of $T_{\mathrm{c,id}}$ when $p\neq 1.5$ (see Appendix C). Also, as a result of this dependence of $T_{\mathrm{c,id}}$ on the potential softening length, we find that the ratio $|T_{\mathrm{p,id}}/T_{\mathrm{p,d}}|$ ends up being roughly independent [^8] In other words, we find that $T_{\mathrm{p,id}}\propto h_{\mathrm{p}}^{-2}$ approximately for $p=0$, which is somewhat coincidentally caused by partial cancellation of $T_{\mathrm{c,id}}\propto h_{\mathrm{p}}^{-1}$ and $T_{\mathrm{L,id}}\propto h_{\mathrm{p}}^{0}$ (they have opposite signs), see Appendix C. of $h_{\mathrm{p}}$ for $p=0$, see Figure 9c.
@@
The small relative contribution of $T_{\mathrm{p,id}}$ to the full planetary torque $T_{\mathrm{p}}$, obvious from the results presented in Figure 9, is typical when the disc is not strongly perturbed by the planet and is smooth around its orbit (a situation realized for $M_{\mathrm{p}}\lesssim M_{\mathrm{th}}$), in which case the power-law $\Sigma_{0}$ profile (7) is appropriate. In this case, most of the torque acting on the planet arises due to the strong direct gravitational coupling at $m\sim h_{\mathrm{p}}^{-1}$ Lindblad resonances located near the planetary orbit (GT80), and $T_{\mathrm{p,id}}$ pales in comparison. As a result, the calculations of the rate of Type I planetary migration (Ward1997) should not be strongly affected when the indirect force due to disc is neglected altogether (Fairbairn2025a).

### v3

@@ 4.3 Deposition torque density
The decay of AMF $F_{J}$ relative to integrated torque $T$ in Figure 7 at large $R$ is caused by the dissipation of the density wave. Our simulations do not have any explicit sources of wave dissipation (e.g. viscosity or radiation), and numerical dissipation is very small, so that wave damping is driven predominantly by nonlinear effects — steepening of the wave profile leading to its shocking and dissipation (Goodman and Rafikov, 2001; Rafikov, 2002a). This mechanism is effective even for low $M_{\mathrm{p}}$, when the nonlinearity is weak. But in a purely linear calculation this process and the associated wave dissipation are absent by default, as Figure 6 demonstrates.
@@
as follows from the wave angular momentum conservation. Non-zero $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ is what ultimately drives disc evolution, leading to phenomena such as gap formation (Rafikov, 2002b; Kanagawa et al., 2015; Cordwell and Rafikov, 2024). Thus, knowing $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ is crucial for understanding evolution of planet-induced disc structures. This motivates us to examine the effect of the IT on the deposition torque density.
@@
We also show $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ obtained in a very different way, namely from a disco simulation with the same parameters but carried out in the barycentric frame, see Figure 8. As indirect forces are absent by default in this frame, this calculation provides an independent check of the effect of the damping density wave on the disc, derived from our astrocentric simulations. The result is shown by the red dot-dashed curve (as a function of barycentric distance $r$, which should not make much difference), and matches the overall shape of $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ quite well. There are some differences, e.g. $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ in the barycentric frame is slightly larger than in the astrocentric frame, but nevertheless the agreement is quite good given the significant differences between the numerical methods used in this comparison.
@@ 4.4 Oscillatory behaviour of the indirect torque
where $H_{\mathrm{p}}$ is the disc scale height at the planetary location. For a disc illustrated in Figure 4, $q=1$ and the distance between the nodes decreases as $R_{k}^{-1/2}$, which is noticeable in the behavior of the green dot-dashed curve in panel (c) of that figure. Equations (26)-(27) are essentially analogous to the periodicity conditions formulated for the torque wiggles in disc-planet coupling (Cimerman et al., 2024) and for the radial oscillations of the gravitational torque density in circumbinary discs (see Eqs. (17), (18) of Cimerman and Rafikov, 2024) since all of these phenomena depend on the global spatial structure of the planet-driven density wave.
Furthermore, the dependence of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ on only the $m=1$ component of $\delta\Sigma$ allows us to explain the intriguing discrepancy between the linear and simulation-based amplitudes of the $\mathrm{d}T/\mathrm{d}R$ oscillations seen in Figure 5. Naively, one might expect the simulation-based $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations (solid) to have lower amplitude than the one predicted by the linear theory (dot-dashed), since in simulations the density wave amplitude decays as $R$ increases due to shock dissipation (see Figure 2a). However, Figure 2c shows that $\delta\Sigma_{1}$ actually increases with $R$, unlike all of the other $\delta\Sigma_{m}$ components. This phenomenon is caused by the nonlinear broadening of the wake as it propagates through the disc, which effectively transfers some power into the $m=1$ mode of the perturbation while the modes with $m>1$ get gradually extinguished by shock damping. This process has been previously described in Cimerman et al. (2024) and is illustrated by their Figure 14. As a result, far from the planet the simulation-based $\delta\Sigma_{1}$ exceeds the linear prediction for $\delta\Sigma_{1}$, explaining the trends in $\mathrm{d}T/\mathrm{d}R$ behavior observed in Figure 5.
@@ 4.5 Divergence of the indirect torque as $R\to\infty$
We now analyze more carefully the origin of the divergence of the indirect torque density $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ as $R\to\infty$, which is clearly visible in Figure 4; this analysis reinforces simple arguments offered in the end of Section 4.1. In principle, the mathematical reason for this divergence is already obvious from equation (14): according to its definition, $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ is proportional to $R^{2}$, which drives the divergence in the first place. However, the actual asymptotic behavior of $\mathrm{d}T_{\mathrm{id}}/\mathrm{d}R$ also depends on the behavior of $\delta\Sigma(R)$, which is something that we can address now.
Equation (25) shows that the amplitude of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations depends on $R$ as $\propto R^{2}|\delta\tilde{\Sigma}_{1}(R)|$, i.e. it increases with $R$ as long as $|\delta\tilde{\Sigma}_{1}(R)|$ decays slower than $R^{-2}$. This is clearly the case in the examples shown in Figures 4c, 6b, 7b, since the oscillation amplitude grows in a divergent fashion as $R\to\infty$. In the linear regime, in the absence of dissipation, one can show (Rafikov, 2002a; Miranda and Rafikov, 2020b) that far in the outer disc, where $\Omega\ll\Omega_{\mathrm{p}}$,
@@
After the wave shocks (which happens earlier for higher $M_{\mathrm{p}}$), the nonlinear wave dissipation causes $|\delta\tilde{\Sigma}_{1}(R)|$ to fall below the linear prediction (28). However, this has a rather limited effect in slowing down the growth of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations [^6] As a result of nonlinear damping, the wave AMF would eventually decay as $F_{J}\propto t^{1/2}$, where the function $t(R)$ is given by the equation (32) of Rafikov (2002a). One can show that in this regime the amplitude of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations would scale as $\propto R^{(11/8)+(7q/16)-(5p/8)}$, not very far from the linear case (28).. Moreover, as Figure 2c illustrates, in low-$M_{\mathrm{p}}$ cases $|\delta\Sigma_{1}|$ may actually increase with $R$ (due to the transfer of power to $m=1$ mode as discussed in the end of Section 4.4) easily enabling the divergence of $\mathrm{d}T_{\mathrm{id}}(R)/\mathrm{d}R$ oscillations in Figure 4c.
@@ 5 Torque on the planet and planetary orbital evolution
We start by emphasizing that in this Section the IT arises because of the reflex motion of the star caused by the gravity of the disc, and not of the planet as was the case in Sections 3 and 4. Taking a cross product of equation (2) with $M_{\mathrm{p}}{\bf R}_{\mathrm{p}}$ one finds that the torque on the planet due to the disc (its only non-zero $z$-component $T_{\mathrm{p}}$) is (Fairbairn and Rafikov, 2025)
$T_{\mathrm{p}}=GM_{\mathrm{p}}\int\limits_{\rm disc}\Sigma({\bf R})\left({\bf R}_{\mathrm{p}}\times{\bf R}\right)_{z}\left(\frac{1}{|{\bf R}_{\mathrm{p}}-{\bf R}|^{3}}-\frac{1}{R^{3}}\right)\mathrm{d}^{2}{\bf R}\,$, (29)
@@
$T_{\mathrm{p,d}}=GM_{\mathrm{p}}\int\limits_{\rm disc}\Sigma({\bf R})\frac{\left({\bf R}_{\mathrm{p}}\times{\bf R}\right)_{z}}{|{\bf R}_{\mathrm{p}}-{\bf R}|^{3}}\,\mathrm{d}^{2}{\bf R}=-\int\limits_{\mathrm{disc}}\frac{\mathrm{d}T_{\mathrm{d}}(R^{\prime})}{\mathrm{d}R}\,\mathrm{d}R^{\prime}$ (30)
@@
$T_{\mathrm{p,id}}=-GM_{\mathrm{p}}\int\limits_{\rm disc}\Sigma({\bf R})\frac{\left({\bf R}_{\mathrm{p}}\times{\bf R}\right)_{z}}{R^{3}}\,\mathrm{d}^{2}{\bf R}.$ (31)
@@
Moreover, this also means that the full angular momentum of the disc-planet system is not conserved in the astrocentric frame. This result, which may seem paradoxical at first sight, is in fact fully in line with the findings of Rafikov (2025), who showed that the angular momentum of a system of gravitating objects — the star, planet and disc, in our case — in a non-inertial frame attached to one of them, is not conserved because of the torque exerted on the system by the indirect force.
With rare exceptions, studies computing the disc torque on the planet accounted only for the direct contribution $T_{\mathrm{p,d}}$, neglecting the indirect part $T_{\mathrm{p,id}}$ altogether. Recently, Fairbairn and Rafikov (2025) explicitly considered $T_{\mathrm{p,id}}$ in their semi-analytical linear study of orbital evolution of low-mass eccentric planets and showed it to be small compared to $T_{\mathrm{p,d}}$. A similar conclusion was subsequently reached in Crida et al. (2025) based on simulations in the context of Type I migration, whereas for more massive (gap-opening) planets they found $T_{\mathrm{p,id}}$ to play a more significant role (see also Regály and Vorobyov, 2017). We now re-assess the relative contribution of the indirect component $T_{\mathrm{p,id}}$ of the disc torque on the planet in our setup using (semi-)analytical linear calculations.
@@ 5.1 Low mass planets
To evaluate the relative contribution of $T_{\mathrm{p,id}}$ to $T_{\mathrm{p}}$, we start by recalling that the one-sided direct planetary torques on the disc are $\sim F_{J,0}$ and similar in magnitude but opposite in sign. The full direct torque given by equation (30) then ends up being (Ward, 1997)
@@
compare with equations (21) and (22). We note that for $R-R_{1}\gtrsim H_{\mathrm{p}}$, outside the $m=1$ OLR at $R_{1}$, $\Sigma_{1}=\delta\Sigma_{1}$ has a form given by the equations (23) and (24). As a result, the integrand in (36) rapidly oscillates in this region, so that this part of the disc provides a small contribution to $T_{\mathrm{p,id}}$. On the other hand, for $R_{1}-R\gtrsim H_{\mathrm{p}}$, inwards of the $m=1$ OLR, $\delta\Sigma_{1}$ decays exponentially as $R$ decreases (see Figure 2b), except at $R=R_{\mathrm{p}}$ where the perturbation is co-orbital with the planet and becomes large. As a result, one expects that the main contributions to the integral in equation (36), as well as in equation (31), come from the immediate vicinity of the $m=1$ OLR, i.e. from $|R-R_{1}|\lesssim H_{\mathrm{p}}$, and from the corotation region $R\approx R_{\mathrm{p}}$. Denoting these contributions as $T_{\mathrm{L,id}}$ (Lindblad) and $T_{\mathrm{c,id}}$ (corotation), respectively, we can thus approximate $T_{\mathrm{p,id}}\approx T_{\mathrm{L,id}}+T_{\mathrm{c,id}}$. We corroborate these expectations in Appendix C, where we also provide an analytical calculation of $T_{\mathrm{L,id}}$.
@@
Fairbairn and Rafikov (2025) studied the torque on the planet from the disc and the resultant planetary orbital evolution using semi-analytical linear theory (Miranda and Rafikov, 2019b; Fairbairn and Rafikov, 2022), as described in Section 2.4. We used their approach to numerically compute both $T_{\mathrm{p,id}}$ and $T_{\mathrm{p,d}}$ and plot their ratio in Figure 10 as a function of different disc parameters: temperature (a) and $\Sigma$ (b) slopes $q$ and $p$, respectively, and the disc aspect ratio $h_{\mathrm{p}}$, for $q=0$ and two values of $p=0$ (c) and $p=1.5$ (d).
@@
To shed light on these trends, we note that in a Keplerian disc with $p=1.5$, one expects the corotation torque to be absent, $T_{\mathrm{c,id}}\to 0$ (Goldreich and Tremaine, 1980). This is discussed in more detail in Appendix C, and is illustrated in Figure 12b. As a result, for this $p$ the indirect torque on the planet is provided only by the Lindblad contribution, $T_{\mathrm{p,id}}\approx T_{\mathrm{L,id}}$, which is known to be independent of $c_{\mathrm{s}}$ or $h_{\mathrm{p}}$ to zeroth order (see Goldreich and Tremaine, 1979, 1980), as underlined by the discussion in Appendix C and our analytical estimate (51). As a result, it follows from equation (35) that $|T_{\mathrm{p,id}}/T_{\mathrm{p,d}}|\propto h_{\mathrm{p}}^{2}$ for $p=1.5$, in agreement with Figure 10d.
@@
On the other hand, in a disc with $p=0$ the indirect corotation torque $T_{\mathrm{c,id}}$ is non-zero and significantly complicates the interpretation of the resultant trends (a similar conclusion has been reached by Crida et al., 2025). We find that in such a disc $T_{\mathrm{c,id}}$ provides a dominant contribution to $T_{\mathrm{p,id}}$ (e.g. see Figure 12a), strongly exceeding that of $T_{\mathrm{L,id}}$. For that reason, $T_{\mathrm{p,id}}$ has a higher amplitude for the $p=0$ disc as compared to the $p=1.5$ disc (the changing relative contributions of $T_{\mathrm{c,id}}$ and $T_{\mathrm{L,id}}$ cause the particular dependence on $p$ shown in Figure 10b). However, it has to be remembered that the magnitude of the corotation torque is a strong function of the potential softening, which itself depends on $h_{\mathrm{p}}$ (our linear calculations here adopt a softening length of $0.3h_{\mathrm{p}}$, following Fairbairn and Rafikov, 2025); hence, there is some ambiguity in the actual amplitude of $T_{\mathrm{c,id}}$ when $p\neq 1.5$ (see Appendix C). Also, as a result of this dependence of $T_{\mathrm{c,id}}$ on the potential softening length, we find that the ratio $|T_{\mathrm{p,id}}/T_{\mathrm{p,d}}|$ ends up being roughly independent [^8] In other words, we find that $T_{\mathrm{p,id}}\propto h_{\mathrm{p}}^{-2}$ approximately for $p=0$, which is somewhat coincidentally caused by partial cancellation of $T_{\mathrm{c,id}}\propto h_{\mathrm{p}}^{-1}$ and $T_{\mathrm{L,id}}\propto h_{\mathrm{p}}^{0}$ (they have opposite signs), see Appendix C. of $h_{\mathrm{p}}$ for $p=0$, see Figure 10c.
@@
The small relative contribution of $T_{\mathrm{p,id}}$ to the full planetary torque $T_{\mathrm{p}}$, obvious from the results presented in Figure 10, is typical when the disc is not strongly perturbed by the planet and is smooth around its orbit (a situation realized for $M_{\mathrm{p}}\lesssim M_{\mathrm{th}}$), in which case the power-law $\Sigma_{0}$ profile (7) is appropriate. In this case, most of the torque acting on the planet arises due to the strong direct gravitational coupling at $m\sim h_{\mathrm{p}}^{-1}$ Lindblad resonances located near the planetary orbit (Goldreich and Tremaine, 1980), and $T_{\mathrm{p,id}}$ pales in comparison. As a result, the calculations of the rate of Type I planetary migration (Ward, 1997) should not be strongly affected when the indirect force due to disc is neglected altogether (Fairbairn and Rafikov, 2025).